OpenStudy (anonymous):

uses the elimination method or the substitution method to solve two linear equations. 5x+3y=17 -5x+2y =3

OpenStudy (amistre64):

these are already primed for elimination 5x+3y=17 -5x+2y =3 ------------ 5y = 20 y = 20/5

OpenStudy (anonymous):

add them together to get 5y =20 so y=4 substituting back 5x+12=17 x=1